High-Low Method The manufacturing costs of Fuld Industries for three months of the year are provided below. January February March Using the high-low method, determine (a) the variable cost per unit and (b) the total fixed cost. Round Total Costs $223,200 302,310 347,200 Production 1,200 units 2,290 3,200 all answers to the nearest whole dollar. a. Variable cost per unit b. Total fixed cost
Answer #1

Variable cost per unit=[Total cost at highest level-Total cost at lowest level]/(Highest level-Lowest level)

=(347200-223200)/(3200-1200)

=$62 per unit

Total fixed cost=347200-(62*3200)=$148,800.